# TIDMAD: Time Series Dataset for Discovering Dark Matter with AI Denoising

### v1

@@ 5 Limitations and applications
As illustrated in Figure 5, the baseline models fail to surpass the results of ABRA Run 3. This is attributed to hardware limitations and changes since the last data run, including (1) replacing the dark matter pickup cylinder with a pickup loop, consequently reducing the geometric coupling to the dark matter signal, and (2) reducing the data-taking period to 24 hours from three months. The decision to implement these changes was driven by the aim to enable ABRA to simultaneously search for dark matter and gravitational waves, thereby enhancing the scientific scope of the experiment. The shortened data-taking duration was necessitated by operational constraints of the dilution refrigerator. While running time series models over the frequency domain data of ABRA Run 3 is not possible, reverting to the previous pickup structure and extending the data-taking period would enable us to generate time series data matching the results of ABRA Run 3.
@@
In Section 4.2, we discussed how to set a dark matter limit using the provided analysis scripts. This script assumes a null result as no $5\sigma$ dark matter candidates were identified in this region of parameter space by ABRA Run 3. Therefore, we assume a null result for this much shorter ($24$ hr) data taking. This assumption enables us to establish upper limits on the coupling parameter $g_{a\gamma\gamma}$ for every mass point. As shown in Figure 5, the ABRA-TIDMAD Raw limit without denoising covers a smaller region than ABRA Run 3 [13].
However, with the denoising algorithm applied to the 24-hour science data, ABRA-TIDMAD could potentially reach beyond the ABRA Run 3 region, where a discovery of dark matter is possible. In this paper, we focused on increasing the experimental sensitivity and setting exclusion limits. A straightforward modification to the interpretation of the TS would unlock the discovery potential of this analysis framework. Future efforts will focus on employing a more extensive dataset and implementing dark matter discovery analysis code. Given that discovering dark matter would be a Nobel Prize-level breakthrough, it is crucial to not only claim discovery but also to convince the scientific community of its validity. If TIDMAD users find any anomalous signals in the science dataset, please contact the authors for further investigation and understandings of systematic uncertainties.
@@
The findings of this study are relevant across a spectrum of physics and computer science applications. Axion dark matter is a subset of wave-like dark matter candidates, therefore the techniques developed in this paper are broadly applicable to all wave-like dark matter experiments. Furthermore, these time series denoising algorithms can be used to search for other wave-like physics signals including gravitational waves. In nuclear and particle physics more broadly, noise reduction is a perennial problem; advancements in denoising can be deployed across a suite of these rare event searches in physics.
While the data released in this paper was tailored to our specific problem statement and benchmarks in physics, these ultra-long time series data hold potential for diverse applications beyond physics. Many long time series datasets, such as stock prices or weather data, exhibit relatively uniform frequency characteristics. As a complement, our data release and subsequent models are built on frequency-rich detector-generated, long time series data and could fill in this gap. If building a foundation model for time series predictions or anomaly detection, our data provide a unique abundance of spectral complexity.
@@ Acknowledgements
J. T. Fry is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. This material is based upon work supported by the National Science Foundation under Grant No. 2110720. We would like to thank San Diego Supercomputer Center (SDSC) and Open Science Data Federation (OSDF) for the data storage space and open access infrastructure. We’d like to thank Frank Wuerthwein for many insightful discussions and offering the OSDF storage space, as well as Fabio Andrijauskas for the technical support. The computation of this project was conducted on SDSC Expanse cluster as well as MIT’s subMIT cluster. We would also like to thank Joshua Foster for his invaluable contributions to the axion analysis.
Appendix A Details and limitations of deep learning model
@@
Since the injected dark matter signal spans two orders of magnitude in frequency, the observed features in the injected time series significantly vary. During training, we noticed that a single deep learning model to denoise the entire dataset would fail to generalize across the different injected frequency ranges. To address this issue, we trained four deep learning models per architecture, each focusing on a specific frequency range: the first covering the low-frequency regime (training/validation files 0-3), the second covering the mid-low regime (training/validation files 4-9), the third covering the mid-high regime (training/validation files 10-14), and the fourth covering the high-frequency regime (training/validation files 15-19). During the benchmark 1 inference, we selected the input validation data corresponding to the frequency range for which each model was trained and averaged the results of the four models. For the benchmark 2 inference, we ran all four models on the science data and selected the highest-performing model, as represented in Figure 5. In future work, we aim to develop a single denoising model generalizable to wide frequency ranges. All trained models are made available in https://drive.google.com/drive/folders/16ORX1b2zo1_lOYYAcRBgddBuYImj0Bxs?usp=share_link.
@@
The transformer model processes the time series data by using an Embedding layer to encode each input, converting 8-bit integers in the range of $(-128,127)$ into a $32$-dimensional vector. Positional encoding is then added to the embedded time series  [16]. This augmented data is fed into a Transformer Encoder with two layers, each containing two heads, $128$ hidden dimensions, and a $0.1$ dropout rate. Finally, the output is passed through a linear layer to produce a $256$-dimensional vector at each time step.
@@
Appendix B Frequentist log-likelihood test statistics
@@
Checklist
@@
    For all authors…

### v2

@@ 5 Limitations and applications
As illustrated in Figure 4.2, the baseline models fail to surpass the results of ABRA-10cm Run 3. This is attributed to hardware limitations and changes since the last data run, including (1) replacing the dark matter pickup cylinder with a pickup loop, consequently reducing the geometric coupling to the dark matter signal, and (2) reducing the data taking period to 24 hours from three months. The decision to implement these changes was driven by the aim to enable ABRA-10cm to simultaneously search for dark matter and gravitational waves, thereby enhancing the scientific scope of the experiment. The shortened data taking duration was necessitated by operational constraints of the dilution refrigerator. Because the signal-to-noise ratio scales as the fourth root of integration time, we can increase ABRA-10cm’s sensitivity to dark matter by increasing the data taking period [18]. Another more efficacious way to increase this signal-to-noise ratio is improving our denoising with ML; a doubling of our noise reduction represents a 16x speed up our data taking time revealing the out-sized return on investment in denoising techniques when compared to increased detector run time.
@@
In Section 4.2, we discussed how to set a dark matter limit using the provided analysis scripts. This script assumes a null result as no $5\sigma$ dark matter candidates were identified in this region of parameter space by ABRA-10cm Run 3. Therefore, we assume a null result for this much shorter ($24$ hr) data taking. This assumption enables us to establish upper limits on the coupling parameter $g_{a\gamma\gamma}$ for every mass point. As shown in Figure 4.2, the ABRA-TIDMAD Raw limit without denoising covers a smaller region than ABRA-10cm Run 3 [25].
However, with the denoising algorithm applied to the 24-hour science data, ABRA-TIDMAD could potentially reach beyond the ABRA-10cm Run 3 region, where a discovery of dark matter is possible. In this paper, we focused on increasing the experimental sensitivity and setting exclusion limits. A straightforward modification to the interpretation of the TS would unlock the discovery potential of this analysis framework. Future efforts will focus on employing a more extensive dataset and implementing dark matter discovery analysis code. Given that discovering dark matter would be a Nobel Prize-level breakthrough, it is crucial to not only claim discovery but also to convince the scientific community of its validity. If TIDMAD users find any anomalous signals in the science dataset, please contact the authors for further investigation and understandings of systematic uncertainties.
@@
Axion dark matter is a specific subset of wave-like dark matter candidates, making the techniques developed in this paper broadly applicable to a wide range of wave-like dark matter experiments. Other axion dark matter experiments, including but not limited to ADMX [7], HAYSTAC [6], and CASPEr [8], also produces long time series data and search for similar peaks in frequency domain; any AI algorithm developed upon TIDMAD can be easily adapted and applied to these experiments. Furthermore, time series denoising algorithms are crucial for extracting wave-like signals in various areas of physics. In astrophysics, gravitational wave searches involve detecting chirp signals with durations on the order of seconds [1], often buried within detector noise. In nuclear physics, denoising can enhance the efficiency of HPGe detectors [5] and bolometer detectors [32]. Advancements in denoising can be deployed across a suite of these physics experiments.
While the data released in this paper was tailored to our specific problem statement and benchmarks in physics, these ultra-long time series datasets have the potential to benefit a wide range of applications beyond physics. Similar to TIDMAD, many other scientific domains involves time series datasets exhibiting relatively uniform frequency characteristics, with the primary analytical task focused on extracting signals from these time series. Examples include pulsar timing from radio observatory data (astronomy)[15], detecting seismic arrivals above background noise (geology)[33], identifying sea surface and near-surface temperature anomalies (climate science)[27], and recognizing atrial fibrillation among noises and other rhythms in short-term ECG recordings (health science)[9]. If a foundation model were to be developed for general time series analysis in science, our frequency-rich, detector-generated, long time series data could provide a uniquely abundant source of spectral complexity.
@@ Acknowledgements
J. T. Fry is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. A. Li’s work is supported by his startup grant. This material is based upon work supported by the National Science Foundation under Grant No. 2110720. We would like to thank San Diego Supercomputer Center (SDSC) and Open Science Data Federation (OSDF) for the data storage space and open access infrastructure. OSDF is funded under National Science Foundation under Grant No. 2030508, 2331480, and 2112167. We’d like to thank Frank Wuerthwein for many insightful discussions and offering the OSDF storage space, as well as Fabio Andrijauskas for the technical support. The computation of this project was conducted on SDSC Expanse cluster as well as MIT’s subMIT cluster. We would also like to thank Joshua Foster for his invaluable contributions to the axion analysis.
Appendix A Dark Matter Signal and Signal Injection
Axion dark matter appears in our detector as a time oscillating current with two free model parameters ($m_{a},g_{a\gamma\gamma}$) described by Equation 1. This signal can be fully derived from axion theory and is extremely well-defined. While a full derivation of this equation is presented in Reference [18], schematically, this theoretical prediction for the dark matter signal in our detection comes from four phenomena -- the omnipresent axion field, axion’s interactions with electromagnetism, the specific geometry of our detector, and the velocity distribution of the dark matter.
Under certain conditions, axions are created in the early universe via the misalignment mechanism [12, 21]. When axions are created they produce a time oscillating omnipresent axion field with an oscillation frequency equal to $m_{a}$ and phase coherence. If axions are dark matter then the abundance of axions and the strength of this axion field are set by our astrophysical observations of dark matter density [11].
The axion field can interact with other physical forces including electromagnetism. If an axion field collides with an electromagnetic field, some of the axions will be transformed into photons. These photons produce a secondary electromagnetic field -- an axion-induced electromagnetic field. In the geometry of our detector, the axion-induced electromagnetic field is read out with a pickup wire which samples the field like a typical radio antenna. This means the time oscillating axion field turns into a time oscillating current in our detector (i.e. our dark matter signal).
If earth were stationary, sitting in a uniform bath of dark matter, then the dark matter signal frequency would be exactly the axion field oscillation frequency (i.e. the axion mass) and completely coherent. However, the earth is moving within the Milky Way galaxy and thus the axion field gains a velocity with respect to earth at about $v_{DM}\sim 220km/s$. Doppler shifting spreads this frequency such that it is distribution around the original field frequency $\Delta f=v_{DM}^{2}f$ [18]. This frequency distribution is six orders of magnitude smaller than the signal frequency, therefore can be treated a coherent, single frequency sine-wave to good approximation. For signal injection, this single frequency approximation is used while for the dark matter analysis, we model the full frequency distribution.
Thus, we have established that the dark matter signal in the ABRA-10cm detector is approximately sine-wave current. The frequency of this oscillating signal is a model parameter meaning if we knew the axion mass, this frequency would be set. However, theoretical models point to a range of possible axion masses, not a singular value. Ideally we would create a dark matter detector that could search the entire range of valid dark matter masses, but experimental constraints such detector size, configuration, and readout electronics preclude this possibility. Instead, experiments must be tailored to search for smaller areas of the axion mass parameter space with ABRA-10cm the detector being specifically designed to target $m_{a}=0.4-10$ neV [19].
The second free signal parameter is the sine wave amplitude which is proportional to $g_{a\gamma\gamma}$, the strength of the axion’s interactions with electromagnetism. Theoretical calculations constrain this parameter to $g_{a\gamma\gamma}=Cm_{a}$ where C=[0.39,0.22]C=[-0.39,0.22] depending on the theory [26, 13]. Theoretically motivated axion couplings can be seen in Figure 4.2 as the gold band in the $m_{a},g_{a\gamma\gamma}$ parameter space. Experimentally, the goal is to detect ever smaller signal amplitudes to reach lower values of $g_{a\gamma\gamma}$.
To inject a fake signal into the hardware, we replicate the signal current, given by Equation 1, with a signal generator. The signal generator is connected to a calibration loop depicted in Figure 1 designed to mimic an axion field incident on the detector. We specifically choose to inject sine waves with frequencies from 1.1 kHz to 4.9 MHz, $m_{a}=[0.005,17]$ neV, to contain the masses our experimental hardware was built to target. The injected fake signals we used all have amplitudes of 50 mV to achieve a reasonable signal-to-noise ratio. While injecting smaller fake signal amplitudes would effectively simulate dark matter candidates with smaller electromagnetic couplings, fake signals smaller than 50 mV are difficult to detect with traditional techniques. Though smaller couplings provide an interesting ML task, our denoising score benchmark is predicated on finding injected signals with traditional techniques and subsequently we did not use smaller fake signal amplitudes. However, we did take data injected with signals ranging from 1.1 kHz to 4.9 MHz at a signal amplitude of 10 mV. While this data was not used in our benchmark creation or model training, it is publicly available (see Datasheet) for an extra challenge.
To summarize, our signal injection scheme involves exciting hardware with a sine wave from a signal generator. We sequentially step through 309 different frequencies from 1.1 kHz to 4.9 MHz to simulated 309 different axion masses in our detector hardware all with an amplitude of 50 mV so that the fake signal is visible above the detector noise floor.
Appendix B Details and limitations of deep learning model
@@
Since the injected dark matter signal spans two orders of magnitude in frequency, the observed features in the injected time series significantly vary. During training, we noticed that a single deep learning model to denoise the entire dataset would fail to generalize across the different injected frequency ranges. To address this issue, we trained four deep learning models per architecture, each focusing on a specific frequency range: the first covering the low-frequency regime (training/validation files 0-3), the second covering the mid-low regime (training/validation files 4-9), the third covering the mid-high regime (training/validation files 10-14), and the fourth covering the high-frequency regime (training/validation files 15-19). During the benchmark 1 inference, we selected the input validation data corresponding to the frequency range for which each model was trained and averaged the results of the four models. For the benchmark 2 inference, we ran all four models on the science data and selected the highest-performing model, as represented in Figure 4.2. In future work, we aim to develop a single denoising model generalizable to wide frequency ranges. All trained models are made available in https://drive.google.com/drive/folders/16ORX1b2zo1_lOYYAcRBgddBuYImj0Bxs?usp=share_link.
@@
The transformer model processes the time series data by using an Embedding layer to encode each input, converting 8-bit integers in the range of (128,127)(-128,127) into a $32$-dimensional vector. Positional encoding is then added to the embedded time series  [31]. This augmented data is fed into a Transformer Encoder with two layers, each containing two heads, $128$ hidden dimensions, and a $0.1$ dropout rate. Finally, the output is passed through a linear layer to produce a $256$-dimensional vector at each time step.
@@
The WaveNet model implements the core dilated causal convolution architecture with 10 residual blocks using exponentially increasing dilation rates (1, 2, 4, 8, 16, ..., 512), kernel size of 12, and 32 residual channels with 64 gate channels for the gated activation mechanism (tanh × sigmoid). Each block processes input through dilated causal convolutions to ensure no future information leakage, applies gated activation, and produces both residual connections (for depth) and skip connections (accumulated across all blocks for the final output). The model uses 32 skip channels and projects the final accumulated skip connections through two 1D convolutions to produce the output shape of (batch, sequence_length, 256) while achieving a large receptive field of over 1000 timesteps through the dilated convolution hierarchy, making it suitable for modeling long-term dependencies in sequential data like audio waveforms.
The simple RNN Seq2Seq model follows the classical encoder-decoder architecture with vocabulary size of 256 and 256 output classes, using 128-dimensional embeddings and 256-dimensional LSTM hidden states across 2 layers with 0.1 dropout. The encoder processes the input sequence (batch, seq_len) through an embedding layer and LSTM to produce final hidden and cell states, which initialize the decoder LSTM that generates output logits of shape (batch, seq_len, num_classes) by processing the same input sequence (a simplified teacher-forcing approach for same-length input/output tasks). The model uses separate embedding and LSTM layers for encoder and decoder, accumulates the decoder’s hidden states across all timesteps, and projects them through a linear layer to produce class distributions for each position, making it suitable for sequence-to-sequence tasks where input and output have identical lengths, such as sequence labeling or token-level classification with 256 possible output categories per position.
Appendix C Frequentist log-likelihood test statistics
@@
Appendix D Current Axion Limits
In Figure 4.2, the TIDMAD dark matter limits, denoised and raw, are presented alongside ABRA-10cm’s previous Run 3 limits and various state-of-the-art axion dark matter experiments and observations. In gold, the theoretically motivated couplings for dark matter candidates are highlighted as discussed in Appendix A. In this appendix, a brief summary of the origins of these limits will be provided.
Light-shining-through-walls:
In this class of experiments, a high-intensity laser is directed towards a solid barrier. While conventional light cannot traverse the barrier, the interaction with the wall may cause a fraction of the light to convert into axions. As a result of their weak interaction with matter, these axions could pass through the wall. On the opposite side, detectors are placed to identify any reconverted light, which would indicate the presence of axions. Light-shining-through-walls experiments must both create axions from photons and detect these axions by converting them back into photons, whereas ABRA-10cm only needs to detect axions, not create them.
Cavity Haloscopes:
This class of experiments generate a strong magnetic field to stimulate axions to convert into microwave photons within a resonant cavity-enclosed space. The resonant cavity is finely tuned to amplify specific frequencies of electromagnetic radiation. Sensitive radio receivers then measure the power within the cavity to identify any potential photon signals indicative of axions. While the conversion mechanism is identical to ABRA-10cm, the resonant cavity only amplifies targeted frequencies, while ABRA-10cm’s readout chain has broadband amplification of axion induced signals. ABRA-10cm, Cavity Haloscopes, and SHAFT are all examples of Haloscopes -- experiments that search for dark matter axions in our galaxy’s dark matter halo.
SHAFT:
The Search for Halo Axions with Ferromagnetic Toroids (SHAFT) experiment is an axion haloscope with a broadband readout, similar to ABRA-10cm in both detection and readout mechanism. There are two main differences between SHAFT and ABRA-10cm (1) SHAFT uses toroidal magnets with ferromagnetic material in the core to convert the axions (2) SHAFT contains a pairs of stacked ferromagnetic toroids each of which has a separate pickup coil and SQUID readout.
Astrophysics:
There are numerous astrophysical processes that would be altered if the axion exists. Broadly, this class of exclusions takes astrophysical observations, calculates how these processes would change if axions exits, and sets limits on possible axion couplings. These limits include the following astrophysical processes. Stellar Cooling. Axions produced in hot astrophysical plasma can transport energy out of stars. This transport of energy critically affects stellar lifetimes, thus observations of stellar energy-loss rates can set limits on axion’s couplings to matter. Photon Flux Large photon fluxes from astrophysical objects like Supernova 1987A traverse the galaxy before being detected terrestrially. Within the galactic magnetic field, some of these supernova photons could be converted into axions. By observing the gamma-ray signals from such events, strong bounds on axions couplings to photons can be derived. Black Hole Superradiance. Light particles, like axions, affect the gravitational waves emitted by black holes through the superradiance mechanism in which axion fields extract energy and angular momentum from the black hole. Observations of stellar black hole spin measurements can therefore constrain allowable axion couplings.
CAST:
The CERN Axion Solar Telescope (CAST) experiment is a prominent axion helioscope. In contrast to haloscopes which search for axions created in the early universe within the dark matter halo surrounding our galaxy, helioscopes search for axions created in our Sun’s heliosphere. CAST uses a strong, movable superconducting magnet to convert axions produced in the core of our Sun into into X-ray photons when aligned with the Sun. CAST is equipped with highly sensitive X-ray detectors at both ends of the magnet, designed to capture these photons. By tracking the Sun and searching for excess X-rays that correlate with solar axions, CAST aims to detect solar axions.
Appendix E Hyperparameters of Moving Average and SG Filter
We conducted additional studies to understand the effects of different hyperparameters of Moving Average and Savitzky–Golay filter algorithms. The Moving Average method has one hyperparameter: window size, while the Savitzky–Golay filter has two hyperparameters: window size and polynomial order.
E.1 Experimental Setup
We designed 60 experimental trials to systematically understand these hyperparameters:
  • Moving Average (10 trials): Window size randomly sampled from the range [10, 5000]
  • Savitzky–Golay filter (50 trials): Window size randomly sampled as odd numbers from [11, 2001], and polynomial order randomly sampled from [2, 20]
For each trial, we executed the corresponding algorithm and calculated both fine and coarse denoising scores using the evaluation procedure described in Section 4. Each tials take about 6 hours on a 8-Core CPU machine with 64Gb memories.
Table 2: Fine and coarse denoising score for Moving Average and Savitzky–Golay filter under different parameters
Algorithms Window Size Polynomial Order Fine Score Coarse Score
None 1.00 1.10
Moving Average 19 - 0.86 0.95
Moving Average 99 - 0.56 0.69
Moving Average 100 - 0.52 0.64
Moving Average 622 - -0.31 -0.47
Moving Average 974 - -0.57 -0.47
Moving Average 1169 - -0.61 -0.72
Moving Average 2632 - -1.07 -0.97
Moving Average 2632 - -1.07 -0.94
Moving Average 2664 - -1.07 -0.94
Moving Average 3040 - -1.23 -1.14
Moving Average 3810 - -1.23 -1.14
SG Filter 19 11 0.95 1.04
SG Filter 171 5 0.58 0.69
SG Filter 945 2 -0.41 -0.31
SG Filter 405 5 0.30 0.42
SG Filter 1245 15 -2.77 -2.35
E.2 Result and Analysis
The results from the 10 Moving Average trials are presented in Table 2. The complete results from all 50 Savitzky–Golay filter trials are displayed in Figure 4.2, with a representative subset also included in Table 2.
Moving Average Performance: The denoising score consistently improves as the window size decreases.
Savitzky–Golay Filter Performance: The algorithm achieves better performance with smaller window sizes and lower polynomial orders. Notably, the best denoising score was achieved in a trial with window size 19 and polynomial order 11. This means the SG fitter fits an 11-dimensional polynomial to every 19 elements in the time series --- representing an extreme overfitting scenario. In other words, SG filter in this trial attempts to memorize and preserve all features in the time series, resulting in a Fine Denoising Score of 0.94, which asymptotically approach but still inferior to the baseline score of 1.0 (when no denoising algorithm is applied).
To validate if moving average has a similar asymptotic behavior, we conducted an additional Moving Average trial with window size 19. This trial, as expected produced the best fine denoising score at 0.85.
Figure 8: Result of all 50 trials of SG filter. The x axis is the window size, the y axis is the polynomial order of SG filter, while the color of each dot represents its fine denoising score.
E.3 Conclusion and Key Finding
Critically, none of the 60 trials surpassed the baseline denoising score of 1.0, where no denoising algorithm is applied. This demonstrates that traditional algorithms are fundamentally inadequate for handling the complex noise characteristics within the TIDMAD dataset. At best, these methods can only asymptotically approximate the raw time series without providing meaningful denoising improvements. These findings underscore the necessity of deep learning-based approaches for effective denoising in the TIDMAD dataset, as traditional signal processing methods prove insufficient for this challenging task.
NeurIPS Paper Checklist
@@ E.3 Conclusion and Key Finding
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the claims made in the paper. The claims in the introduction, abstract, and main paper are the same; we introduce a new, segmented dataset from a physics detector, a denoising score for direct model benchmarking, and a complete dark matter analysis framework for science legible benchmarking.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Included in Section 5 is a discussion of the limitation of this work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section 1 and Section A, we discuss the physics theory and assumptions for our result. Additionally, for the evaluation metric, we provide a set of assumptions for the benchmark construction in Section 4.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We publicly provide all of the data and code necessary to reproduce all experimental results of the paper. Please use the link in the abstract to access all necessary materials for reproduction.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example